OpenStudy (anonymous):

Plutonium-240 decays according to the function \[Q(t)=Q _{o}e ^{-kt}\] where Q represents the quantity remaining after t years and k is the decay constant, 0.00011... How long will it take 36 grams of plutonium-240 to decay to 12 grams? A. 18,900 years B. 1.44 years C. 9,990 years D. 2,100 years

OpenStudy (anonymous):

so Qt = 12 Qo=36

OpenStudy (anonymous):

use the rule of logarithms. take natural log of both sides

OpenStudy (anonymous):

how exactly do i do that? like what would the equation look like?

OpenStudy (anonymous):

right now the equation is: 12=36*e^(-0.00011*t)

OpenStudy (anonymous):

so if you take the natural log of both sides, you get

OpenStudy (anonymous):

ln(12/36)=ln(e^0.00011*t)

OpenStudy (anonymous):

and ln(e^x) = x

OpenStudy (anonymous):

so you get ln(12/36)=0.00011*t

OpenStudy (anonymous):

so t = ln(12/36) / 0.00011

OpenStudy (anonymous):

make sense?

OpenStudy (anonymous):

well that makes sense, but for the answer i got no solution, what are your thoughts?

OpenStudy (anonymous):

the answer should be C per the equation i gave

OpenStudy (anonymous):

also k is negative

OpenStudy (anonymous):

hm, I'll double check my work, thank you :)

OpenStudy (anonymous):

if it decays, then k is negative

OpenStudy (anonymous):

Answer should be a